OpenStudy (anonymous):

∫sinxcosx from pi/6 to pi/3

myininaya (myininaya):

let u=sinx du=cosx dx \[\int\limits_{}^{}u du=\frac{u^{1+1}}{1+1}+C=\frac{u^2}{2}+C=\frac{(sinx)^2}{2}+C\]

myininaya (myininaya):

plug in limits

OpenStudy (anonymous):

i got that then i plugged in the values and got 0

OpenStudy (anonymous):

actually i got sin2x/2

myininaya (myininaya):

\[\frac{(\sin(\frac{\pi}{3}))^2}{2}-\frac{(\sin(\frac{\pi}{6}))^2}{2}\] =\[\frac{(\frac{\sqrt{3}}{2})^2}{2}-\frac{(\frac{1}{2})^2}{2}\] =\[\frac{1}{2}*\frac{3}{4}-\frac{1}{2}*\frac{1}{4}=\frac{3}{8}-\frac{1}{8}=\frac{2}{8}=\frac{1}{4}\]

OpenStudy (anonymous):

can u explain to me how you integrated it

myininaya (myininaya):

do you see my substitution above?

myininaya (myininaya):

i let u=sinx so du=cosxdx

myininaya (myininaya):

you could also do : remember sin(2x)=sin(x+x)=sinxcosx+sinxcosx=2sinxcosx \[\int\limits_{}^{}\frac{1}{2}*\sin(2x)dx\]

OpenStudy (anonymous):

well first the question asked to show that sin2x=2sinxcosx i proved it using Sin(A+B).

OpenStudy (anonymous):

yeah i got that answer

OpenStudy (anonymous):

but once i put in the values it ends up being 0

myininaya (myininaya):

let u=2x du=2 dx \[\frac{1}{2}\int\limits_{}^{}\frac{1}{2}*2\sin(2x)dx=\frac{1}{2}*\frac{1}{2}\int\limits_{}^{}sinudu=\frac{1}{4}*(-cosu)+C\] =\[\frac{-1}{4}\cos(2x)+C\]

OpenStudy (anonymous):

oh lol i forgot to integrate 1/2 (sin2x)

myininaya (myininaya):

\[\frac{-1}{4}(\cos(2*\frac{\pi}{3})-\cos(\frac{2*\pi}{6}))=\frac{-1}{4}(\cos(\frac{2\pi}{3})-\cos(\frac{\pi}{3}))\] \[=\frac{-1}{4}(\frac{-1}{2}-\frac{1}{2})=\frac{-1}{4}(-1)=\frac{1}{4}\]

myininaya (myininaya):

which way you like better?

myininaya (myininaya):

brb

OpenStudy (anonymous):

i think the second one lol

OpenStudy (anonymous):

whenever substitution is invovled i always manage to stuff it up